ANSWER ALL OF THE QUESTIONS AND CLEARLY SHOW YOUR WORK PLZZZZZZZZZZZZZZZZZZZZZZZZZZZZ

A medication is ordered by the physician 2 g in 250 mL NS to be administered by continuous IV around the clock fo 24 hours. TThe label instructions for KEFUROX specifies that when 77 mL of sterile water is added to a 7.5 g vial, the re 80 mL solutionThe recommended dose for a pediatric antibiotic is 40 mg/m2. If the package insert instruction specifies that when 7 mL of diPenicillin G sodium is available in the hospital pharmacy in vials containing 5,000,000 units of drug. The directions on theA sterile solution request form is received for a large volume parenteral. The infusion rate is 125 mL/h. 500 mL bags are req8 hours. A nurse is hanging a 2% lidocaine solution for a patient. If the dose is 4 mg/min, a 100 mL bag last for Round answe
Answer #1

2g in 250 ml administered over 24 hours

Rate of infusion is 100ml per hour

So for 24 hours we need 24*100 = 2400ml

We have 2g in 250 so it will be 2400/250 = 9.6

Approximately 10

77ml of sterile water to 7.5 gm vial

80ml of cefuroxime concentration is 468.75mg per teaspoon

One teaspoon = 5ml so the concentration of cefuroxime is 468.75mg/5 = 93.75 mg

Weight of child 50lb ie 22.6796kg ( 1lbs = 0.453592kg)

Recommended dose 6mg/kg

So, for the child it will be 6*22.68= 136.08mg

Recommended dose of antibiotic 40mg/m2

7ml of diluent produce 10ml of suspension with 250mg per teaspoon

That is 250/5 = 50mg/ml

Recommended dose is 15 drops in 20ml/drops

So the recommended dose in ml will 15/20 = 0.75 ml

Surface area of 0.5m2 so the dose will be 40mg* 0.5 = 20mg

So the child getting 20 mg in .75 ml so the dilution is 20/0.75 = 26.67mg/ml

The diluent already has 3ml along with 2ml can give approximate 25mg/ml of dilution as per the dose

Answer should be 2

penicillin g sodium 50 lakh units

23 ml produce 2lakhs per ml

That means 23*2 plus 4 ml additional

Need to prepare 1lakh unit per ml

So we have to ad 46 that is 23*2 to make it one lakh per ml

Solution at 125ml per hour for 24 hours

So total solution needed is 125*24= 3000ml

Each bag contains 500ml

So total bags needed = 3000/500 = 6

2% lidocaine solution

That is 2g in 100ml

So, 2000mg in 100ml

Dose is 4mg/min

That is 240mg/hour

So 100 ml should be given in 2000/240 hours = 8.33

Answer is 8